Alvin is 30 years older than his daughter ezra. 20 years from now he will be 4 years more than twice than ezra . What are their present age ?
Answer by fractalier(6550)

Call their ages NOW, a and e.
Then we have
a = e + 30 and
a + 20 = 2(e + 20) + 4
Now substitute in for a in the second equation and get
e + 30 + 20 = 2(e + 20) + 4
e + 50 = 2e + 40 + 4
e + 50 = 2e + 44
e + 6 = 2e
e = 6
a is then 36.